OpenStudy (anonymous):

Express the following equation in slope-intercept form: 5x - 6y = -12

OpenStudy (anonymous):

y=5/6x+2

OpenStudy (anonymous):

slope intercept form is the y=mx+b form you see all the time. you have to get y by itself, so you subtract 5x from both sides and you'll get -6y=-12-5x and then divide both sides by -6 to get y alone. the final equation should be y=5/6x+2